# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v1

@@ F.3 Proofs of Lemmas for Theorem 3
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}.$ (F.2)
@@
$\Leftrightarrow\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})+\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}.$ (F.3)
@@
where the first inequality holds by (F.3), and the last inequality holds since $f(x)=\frac{x}{v_{0}+x}$ is an increasing function.
Case 2. $i^{\star}\notin S_{t}$.
Let us return to (F.2). Since $\frac{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\geq 1$ for any $S\in\mathcal{S}$, we have
@@
F.3.2 Proof of Lemma F.2
Proof of Lemma F.2.
We prove by contradiction. Assume that there exists $i\in S^{\star}$ such that $r_{i}<R(S^{\star},\mathbf{w}_{V})$. Then, removing the item $i$ from assortment $S^{\star}$ yields higher expected revenue. This contradicts the optimality of $S^{\star}$. Thus, we have
$r_{i}\geq R(S^{\star},\mathbf{w}_{V}),\quad\forall i\in S^{\star}.$
This concludes the proof. ∎
Appendix G Proofs of Theorem 4
@@
We begin the proof by introducing additional useful lemmas. Lemma G.1 shows that $\tilde{R}_{t}(S_{t})$, defined in (6), is an upper bound of the true expected revenue of the optimal assortment, $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})$.
Lemma G.1 (Lemma 4 in Oh and Iyengar 22).
Let $\tilde{R}_{t}(S)=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{ti})}$. And suppose $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$. If for every item $i\in S_{t}^{\star}$, $\alpha_{ti}\geq x_{ti}^{\top}\mathbf{w}^{\star}$, then for all $t\geq 1$, the following inequalities hold:
@@
Note that Lemma G.1 does not claim that the expected revenue is a monotone function in general. Instead, it specifically states that the value of the expected revenue, when associated with the optimal assortment, increases with an increase in the MNL parameters [5, 22].
Furthermore, we provide an elliptical potential Lemma G.2 for the centralized context vectors $\tilde{x}_{ti}$.
Lemma G.2.
Let $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$, where $\mathcal{G}_{s}(\mathbf{w})=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w})p_{s}(j|S_{s},\mathbf{w})x_{si}x_{sj}^{\top}$. Define $\tilde{x}_{si}=x_{si}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}_{s+1})}[x_{sj}]$. Suppose $\lambda\geq 1$. Then the following statements hold true:
@@ G.1 Proof of Theorem 4
By Lemma G.1, we can bound the regret as follows:
$\sum_{t=1}^{T}R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})$
Now, we define $\tilde{Q}:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})\in\mathbb{R}^{|S_{t}|}$, $\tilde{Q}(\mathbf{u})=\sum_{i=1}^{|S_{t}|}\frac{\exp(u_{i})r_{ti}}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Let $S_{t}=\{i_{1},\dots,i_{|S_{t}|}\}$. Moreover, for all $t\geq 1$, let $\mathbf{u}_{t}=(u_{ti_{1}},\dots u_{ti_{|S_{t}|}})^{\top}=(\alpha_{ti_{1}},\dots,\alpha_{ti_{{|S_{t}|}}})^{\top}$ and $\mathbf{u}^{\star}_{t}=(u_{ti_{1}}^{\star},\dots u_{ti_{|S_{t}|}}^{\star})^{\top}=(x_{ti_{1}}^{\top}\mathbf{w}^{\star},\dots,x_{ti_{{|S_{t}|}}}^{\top}\mathbf{w}^{\star})^{\top}$. Then, by applying a second order Taylor expansion, we obtain
$\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})$ $=\sum_{t=1}^{T}\tilde{Q}(\mathbf{u}_{t})-\tilde{Q}(\mathbf{u}^{\star}_{t})$
@@
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}\frac{\exp(x_{ti}^{\top}\mathbf{w}^{\star})r_{ti}}{v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star})}(u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}\frac{\exp(x_{tj}^{\top}\mathbf{w}^{\star})r_{tj}\sum_{i\in S_{t}}\exp(x_{tj}^{\top}\mathbf{w}^{\star})}{(v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star}))^{2}}(u_{ti}-u_{ti}^{\star})$
@@
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left((u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{tj}-u_{tj}^{\star})\right)$
$\leq\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left((u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{tj}-u_{tj}^{\star})\right),$
where in the first inequality, we denote $S_{t}^{+}$ as the subset of items in $S_{t}$ such that the term $(u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{tj}-u_{tj}^{\star})$ greater than or equal to zero, and the last inequality holds since $r_{ti}\leq 1$. Let $x_{t0}=\mathbf{0}$. Then, we can further bound the right-hand side as follows:
$\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left((u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})(u_{tj}-u_{tj}^{\star})\right)$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}\cup\{0\}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
@@
where the first inequality holds by Lemma D.1, the second inequality due to Jensen’s inequality, and the second-to-last inequality holds due to the fact that $\|a\|=\|a-b+b\|\leq\|a-b\|+\|b\|$ for any vectors $a,b\in\mathbb{R}^{d}$.
For simplicity, we denote $\mathbb{E}_{\mathbf{w}}[x_{tj}]:=\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w})}[x_{tj}]$. Let $\bar{x}_{ti}=x_{ti}-\mathbb{E}_{\mathbf{w}^{\star}}[x_{tj}]$ and $\tilde{x}_{ti}=x_{ti}-\mathbb{E}_{\mathbf{w}_{t+1}}[x_{tj}]$. Then, we have
@@
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}},$ (G.1)
where the first inequality holds by the triangle inequality. Now, we bound the terms on the right-hand side of (G.1) individually. For the first term, we have
@@
$=\beta_{t+1}(\delta)\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2}+\beta_{t+1}(\delta)\left(\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}\right)^{2}$
@@
where the fourth inequality holds due to Jensen’s inequality and the second-to-last inequality holds by Lemma 1. Hence, we have
@@
$\leq\frac{4}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),$ (G.2)
where the last inequality holds by Lemma D.2. Using similar reasoning, we can bound the second term of (G.1). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
$\leq\beta_{t+1}(\delta)\max_{i\in S_{t}}\frac{\|x_{ti}\|_{H_{t}^{-1}}^{2}+\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}}{2}+\beta_{t+1}(\delta)\frac{\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}+\max_{k\in S_{t}}\|x_{tk}\|_{H_{t}^{-1}}^{2}}{2}$
@@
where the second inequality holds since $(\max_{i}a_{i})^{2}=\max_{i}a_{i}^{2}$ for any $a_{i}\geq 0$. Thus, by Lemma D.2 (or Lemma G.2), we get
@@
$\leq 2\beta_{t+1}(\delta)\max\left\{\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right\}\leq\frac{4}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),,$ (G.3)
Finally, we can bound the third term of (G.1). By the Cauchy-Schwarz inequality, we have
@@
$\leq\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)},$ (G.4)
where the last inequality holds by Lemma G.2. Plugging (G.2), (G.3), and (G.4) into (G.1), we get
@@
$\sum_{t=1}^{T}\nabla\tilde{Q}(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})\leq 2\beta_{T}(\delta)\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)}+\frac{16}{\kappa}\beta_{T}(\delta)\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),$ (G.5)
Now, we bound the term (D). Define $Q:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})\in\mathbb{R}^{|S_{t}|}$, $Q(\mathbf{u})=\sum_{i=1}^{|S_{t}|}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Let $p_{i}(\bar{\mathbf{u}}_{t})=\frac{\exp(\bar{u}_{ti})}{v_{0}+\sum_{k=1}^{K}\exp(\bar{u}_{tk})}$. By following the reasoning presented in Section D.1, we have
@@
$\leq 10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},$ (G.6)
where the first inequality holds since $r_{ti}\in[0,1]$. Combining (G.5) and  (G.6), we derive that

### v6

@@ G.3 Proofs of Lemmas for Theorem 3
$\gamma=\min_{S\in\mathcal{S}}\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{v_{0}+\min_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}.$ (G.2)
@@
$\Leftrightarrow\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})+\sum_{i\in S\setminus\{i^{\star}\}}\exp(x_{i}^{\top}\mathbf{w}_{V})\gamma}{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\leq\frac{\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}{v_{0}+\exp(x_{i^{\star}}^{\top}\mathbf{w}_{V})}.$ (G.3)
@@
where the first inequality holds by (G.3), and the last inequality holds since $f(x)=\frac{x}{v_{0}+x}$ is an increasing function.
Case 2. $i^{\star}\notin S$.
Let us return to (G.2). Since $\frac{v_{0}+\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}{\sum_{i\in S}\exp(x_{i}^{\top}\mathbf{w}_{V})}\geq 1$ for any $S\in\mathcal{S}$, we have
@@
Appendix H Proofs of Theorem 4
@@
We begin the proof by introducing additional useful lemmas. Lemma H.1 shows that $\tilde{R}_{t}(S_{t})$, defined in (6), is an upper bound of the true expected revenue of the optimal assortment, $R_{t}(S_{t}^{\star},\mathbf{w}^{\star})$.
Lemma H.1 (Lemma 4 in Oh and Iyengar 40).
Let $\tilde{R}_{t}(S)=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$. And suppose $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$. If for every item $i\in S_{t}^{\star}$, $\alpha_{ti}\geq x_{ti}^{\top}\mathbf{w}^{\star}$, then for all $t\geq 1$, the following inequalities hold:
@@
Note that Lemma H.1 does not claim that the expected revenue is a monotone function in general. Instead, it specifically states that the value of the expected revenue, when associated with the optimal assortment, increases with an increase in the MNL parameters [7, 40].
Lemma H.2 shows that $\tilde{R}_{t}(S_{t})$ increases as the utilities of items in $S_{t}$ increase.
Lemma H.2.
Let $\tilde{R}_{t}(S)=\frac{\sum_{i\in S}\exp(\alpha_{ti})r_{ti}}{v_{0}+\sum_{j\in S}\exp(\alpha_{tj})}$ and $S_{t}=\mathop{\mathrm{argmax}}_{S\in\mathcal{S}}\tilde{R}_{t}(S)$. Assume $\alpha^{\prime}_{ti}\geq\alpha_{ti}\geq 0$ for all $i\in[N]$. Then, we have
$\tilde{R}_{t}(S_{t})\leq\frac{\sum_{i\in S_{t}}\exp(\alpha^{\prime}_{ti})r_{ti}}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha^{\prime}_{tj})}.$
Furthermore, we provide a novel elliptical potential Lemma H.3 for the centralized context vectors $\tilde{x}_{ti}$.
Lemma H.3.
Let $H_{t}=\lambda\mathbf{I}_{d}+\sum_{s=1}^{t-1}\mathcal{G}_{s}(\mathbf{w}_{s+1})$, where $\mathcal{G}_{s}(\mathbf{w})=\sum_{i\in S_{s}}p_{s}(i|S_{s},\mathbf{w})x_{si}x_{si}^{\top}-\sum_{i\in S_{s}}\sum_{j\in S_{s}}p_{s}(i|S_{s},\mathbf{w})p_{s}(j|S_{s},\mathbf{w})x_{si}x_{sj}^{\top}$. Define $\tilde{x}_{si}=x_{si}-\mathbb{E}_{j\sim p_{s}(\cdot|S_{s},\mathbf{w}_{s+1})}[x_{sj}]$. Suppose $\lambda\geq 2$. Then the following statements hold true:
@@ H.1 Proof of Theorem 4
Let $\alpha^{\prime}_{ti}=x_{ti}^{\top}\mathbf{w}^{\star}+2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}$. If $\mathbf{w}^{\star}\in\mathcal{C}_{t}(\delta)$, then, by Lemma E.1, we have
$\alpha_{ti}\leq x_{ti}^{\top}\mathbf{w}^{\star}+2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}=\alpha^{\prime}_{ti}.$
We denote $\accentset{\approx}{R}_{t}(S_{t})=\frac{\sum_{i\in S_{t}}\exp(\alpha^{\prime}_{ti})r_{ti}}{v_{0}+\sum_{j\in S_{t}}\exp(\alpha^{\prime}_{tj})}$. Then, we can bound the regret as follows:
$\sum_{t=1}^{T}R_{t}(S_{t}^{\star},\mathbf{w}^{\star})-R_{t}(S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\tilde{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\accentset{\approx}{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star}),$
where the first inequality holds by Lemma H.1 and the last inequality holds by Lemma H.2.
Now, we define $\tilde{Q}:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})^{\top}\in\mathbb{R}^{|S_{t}|}$, $\tilde{Q}(\mathbf{u})=\sum_{k=1}^{|S_{t}|}\frac{\exp(u_{k})r_{ti_{k}}}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Let $S_{t}=\{i_{1},\dots,i_{|S_{t}|}\}$. Moreover, for all $t\geq 1$, let $\mathbf{u}_{t}=(u_{ti_{1}},\dots u_{ti_{|S_{t}|}})^{\top}=(\alpha^{\prime}_{ti_{1}},\dots,\alpha^{\prime}_{ti_{{|S_{t}|}}})^{\top}$ and $\mathbf{u}^{\star}_{t}=(u_{ti_{1}}^{\star},\dots u_{ti_{|S_{t}|}}^{\star})^{\top}=(x_{ti_{1}}^{\top}\mathbf{w}^{\star},\dots,x_{ti_{{|S_{t}|}}}^{\top}\mathbf{w}^{\star})^{\top}$. Then, by applying a second order Taylor expansion, we obtain
$\sum_{t=1}^{T}\accentset{\approx}{R}_{t}(S_{t})-R_{t}(S_{t},\mathbf{w}^{\star})$ $=\sum_{t=1}^{T}\tilde{Q}(\mathbf{u}_{t})-\tilde{Q}(\mathbf{u}^{\star}_{t})$
@@
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}\frac{\exp(x_{ti}^{\top}\mathbf{w}^{\star})r_{ti}}{v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star})}(u_{ti}-u_{ti}^{\star})-\sum_{j\in S_{t}}\frac{\exp(x_{tj}^{\top}\mathbf{w}^{\star})r_{tj}\sum_{i\in S_{t}}\exp(x_{ti}^{\top}\mathbf{w}^{\star})}{(v_{0}+\sum_{k\in S_{t}}\exp(x_{tk}^{\top}\mathbf{w}^{\star}))^{2}}(u_{ti}-u_{ti}^{\star})$
@@
$=\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(2\beta_{t}(\delta)\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})2\beta_{t}(\delta)\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right).$
Let $x_{t0}=\mathbf{0}$. Then, we can further bound the right-hand side as follows:
$2\sum_{t=1}^{T}\beta_{t}(\delta)$ $\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\sum_{j\in S_{t}\cup\{0\}}p_{t}(j|S_{t},\mathbf{w}^{\star})\|x_{tj}\|_{H_{t}^{-1}}\right)$
$=2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})r_{ti}\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
$\leq 2\sum_{t=1}^{T}\beta_{t}(\delta)\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
$\leq 2\beta_{T}(\delta)\sum_{t=1}^{T}\sum_{i\in S_{t}^{+}}p_{t}(i|S_{t},\mathbf{w}^{\star})\left(\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\right)$
@@
where, in the first inequality, we define $S_{t}^{+}\subseteq S_{t}$ as the subset of items in $S_{t}$ such that the term $\|x_{ti}\|_{H_{t}^{-1}}-\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w}^{\star})}\left[\|x_{tj}\|_{H_{t}^{-1}}\right]\geq 0$ and $r_{ti}\in[0,1]$, the second inequality holds because $\beta_{1}(\delta)\leq\dots\leq\beta_{T}(\delta)$, the third inequality holds due to Jensen’s inequality, and the second-to-last inequality holds due to the fact that $\|\mathbf{a}\|=\|\mathbf{a}-\mathbf{b}+\mathbf{b}\|\leq\|\mathbf{a}-\mathbf{b}\|+\|\mathbf{b}\|$ for any vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{d}$.
For simplicity, we denote $\mathbb{E}_{\mathbf{w}}[x_{tj}]=\mathbb{E}_{j\sim p_{t}(\cdot|S_{t},\mathbf{w})}[x_{tj}]$. Let $\bar{x}_{ti}=x_{ti}-\mathbb{E}_{\mathbf{w}^{\star}}[x_{tj}]$ and $\tilde{x}_{ti}=x_{ti}-\mathbb{E}_{\mathbf{w}_{t+1}}[x_{tj}]$. Then, we have
@@
$+\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}_{t+1})\|\tilde{x}_{ti}\|_{H_{t}^{-1}},$ (H.1)
where the inequality holds by the triangle inequality. Now, we bound the terms on the right-hand side of (H.1) individually. For the first term, we have
@@
$\leq\beta_{t+1}(\delta)\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}^{2}+\beta_{t+1}(\delta)\left(\sum_{j\in S_{t}}p_{t}(j|S_{t},\bm{\xi}_{t})\left\|x_{tj}\right\|_{H_{t}^{-1}}\right)^{2}$
@@
where the fourth inequality holds by Lemma 1 and the second-to-last inequality holds due to Jensen’s inequality. Hence, we have
@@
$\leq\frac{4}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),$ (H.2)
where the last inequality holds by Lemma E.2. Using similar reasoning, we can bound the second term of (H.1). By the mean value theorem, there exists $\bm{\xi}^{\prime}_{t}=(1-c^{\prime})\mathbf{w}^{\star}+c^{\prime}\mathbf{w}_{t+1}$ for some $c^{\prime}\in(0,1)$ such that
@@
$=\beta_{t+1}(\delta)\max_{i\in S_{t}}\frac{\|x_{ti}\|_{H_{t}^{-1}}^{2}+\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}}{2}+\beta_{t+1}(\delta)\frac{\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}+\max_{k\in S_{t}}\|x_{tk}\|_{H_{t}^{-1}}^{2}}{2}$
@@
where the equality holds since $(\max_{i}a_{i})^{2}=\max_{i}a_{i}^{2}$ for any $a_{i}\geq 0$. Thus, by Lemma H.3 (or Lemma E.2), we get
@@
$\leq 2\beta_{t+1}(\delta)\max\left\{\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2},\max_{i\in S_{t}}\|\tilde{x}_{ti}\|_{H_{t}^{-1}}^{2}\right\}\leq\frac{4}{\kappa}\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),,$ (H.3)
Finally, we bound the third term of (H.1). By the Cauchy-Schwarz inequality, we have
@@
$\leq\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)},$ (H.4)
where the last inequality holds by Lemma H.3. Plugging (H.2), (H.3), and (H.4) into (H.1), we get
@@
$\sum_{t=1}^{T}\nabla\tilde{Q}(\mathbf{u}_{t}^{\star})^{\top}(\mathbf{u}_{t}-\mathbf{u}^{\star}_{t})\leq 2\beta_{T}(\delta)\sqrt{T}\sqrt{2d\log\left(1+\frac{T}{d\lambda}\right)}+\frac{16}{\kappa}\beta_{T}(\delta)\beta_{T+1}(\delta)d\log\left(1+\frac{T}{d\lambda}\right),$ (H.5)
Now, we bound the term (D). Define $Q:\mathbb{R}^{|S_{t}|}\rightarrow\mathbb{R}$, such that for all $\mathbf{u}=(u_{1},\dots,u_{|S_{t}|})\in\mathbb{R}^{|S_{t}|}$, $Q(\mathbf{u})=\sum_{i=1}^{|S_{t}|}\frac{\exp(u_{i})}{v_{0}+\sum_{j=1}^{|S_{t}|}\exp(u_{j})}$. Then, we have $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$ since $r_{ti}\in[0,1]$. By following the similar reasoning from the equation (E.3) to (E.5) in Section E.1, we have
@@
$\leq 10\beta_{T}(\delta)^{2}\sum_{t=1}^{T}\max_{i\in S_{t}}\|x_{ti}\|_{H_{t}^{-1}}^{2}.$ (H.6)
where the first inequality holds because $\left|\frac{\partial^{2}\tilde{Q}}{\partial i\partial j}\right|\leq\left|\frac{\partial^{2}Q}{\partial i\partial j}\right|$. Combining (H.5) and  (H.6), we derive that